celticcat:

Solve for x ; x^3 = 9

happyday09:

x=3 I think

happyday09:

but don't trust that, I'm no good at math

Sprout:

So if x^3=9 is your equation, what do you think you need to do first?

Sprout:

Well it's quite simple. In order to get the answer and x by itself you must take the cube root of the x^3 and 9. When you do that, x=2.08

Sprout:

Also, this question belongs under mathematics, not OpenStudy Memorial. 😁

gn:

@sprout wrote:
Also, this question belongs under mathematics, not OpenStudy Memorial. 😁
It is under mathematics though..

Spectrum:

@gn wrote:
@sprout wrote:
Also, this question belongs under mathematics, not OpenStudy Memorial. 😁
It is under mathematics though..
A moderator had switched the subject it was under.

Spectrum:

@gdub08 wrote:
3
That is the square root of 9...

Gdub08:

@spectrum wrote:
Dude the ^ means exponent
wait so x cubed?

Spectrum:

Yes. x^3 is x cubed

Gdub08:

@spectrum wrote:
Yes. x^3 is x cubed
ok hang on

Gdub08:

wouldn't it give u a decimal?

Gdub08:

2.0888?...

Spectrum:

@sprout wrote:
Well it's quite simple. In order to get the answer and x by itself you must take the cube root of the x^3 and 9. When you do that, x=2.08
Yes, as sprout said. This is the answer rounded to the nearest 100th.

Gdub08:

@spectrum wrote:
@sprout wrote:
Well it's quite simple. In order to get the answer and x by itself you must take the cube root of the x^3 and 9. When you do that, x=2.08
Yes, as sprout said. This is the answer rounded to the nearest 100th.
I'm so confused. It just gives you a very long decimal